# Kolmogorov-Arnold Energy Models: Fast and Interpretable Generative Modeling

### v8

@@ 2.1 Univariate energy-based prior
Here, $F^{-1}_{\pi_{q,p}}:[0,1]\to\mathcal{Z}_{q,p}$ is the generalized inverse CDF associated with a parameterized latent measure, $\pi_{q,p}$ in the latent space, $\mathcal{Z}_{q,p}\subset\mathbb{R}$. We have therefore interpreted $u_{p}\sim\text{Unif}(0,1)$ in Eq. 1 as a uniform random variable in the probability space, $\left([0,1],\mathcal{B}([0,1]),\mu\right)$, where $\mathcal{B}\left([0,1]\right)$ is the Borel $\sigma$-algebra on $[0,1]$ and $\mu$ is the uniform probability measure on $[0,1]$.
This is valid under standard regularity assumptions on $F$, since it is a measurable, monotone, increasing function, and thus $F_{\pi_{q,p}}^{-1}(u)=\inf\left\{z\in\mathcal{Z}_{q,p}:F_{\pi_{q,p}}(z)\geq u\right\}$ is a measurable pushforward (or Markov kernel) between the two probability spaces:
$\left([0,1],\mathcal{B}\left([0,1]\right),\mu\right)\;\underset{F_{\pi_{q,p}}^{-1}}{\longrightarrow}\;\left(\mathcal{Z}_{q,p},\mathcal{B}\left(\mathcal{Z}_{q,p}\right),\pi_{q,p}\right),$ (3)
We parameterize the latent measure $\pi_{q,p}$ via exponential tilting of a base measure $\pi_{0}$:
@@
where $f_{q,p}(z)$ is a learned energy function and $Z_{q,p}$ is the finite normalizing constant (partition function) that ensures $\pi_{q,p}$ induces a valid probability density, $p_{q,p}$, for each latent feature.
The inspiration for this exponential tilting form can be traced to the energy-based prior introduced by pang2020learninglatentspaceenergybased, adapted to the univariate case. The base prior $\pi_{0}$ serves as a reference that guides the form of the learned $\pi_{q,p}$, allowing for flexible reweighting via the energy function $f_{q,p}$. For $\pi_{q,p}$ to define a valid density, which we denote $p_{q,p}$, we require exponential integrability:
@@
$\text{such that}\quad Z_{q,p}$ $=\int_{\mathcal{Z}_{q,p}}\exp(f_{q,p}(z))\,\pi_{0}(dz)<\infty.$ (5)
In practice, $f_{q,p}$ can be parameterized in various ways, made possible by the recent development of Kolmogorov-Arnold Networks (KANs) (liu2024kankolmogorovarnoldnetworks). For instance, smoothing splines, such as those used by KANs or deep spline networks, (9264754). Splines are typically defined on bounded domains, allowing the support of $\pi_{q,p}$ to be explicitly controlled by the user using the KAN grids. These grids must be periodically updated during training according to input activations, following the implementation described by liu2024kankolmogorovarnoldnetworks.
The partition function $Z_{q,p}$ is often intractable in energy-based models. However, in this univariate setting, where the energy function is smooth and defined on a bounded domain $\left[z_{\text{min}},z_{\text{max}}\right]$, it can be accurately approximated using Gauss-Kronrod quadrature, (gauss-quad):
$Z_{q,p}$ $=\int_{z_{\text{min}}}^{z_{\text{max}}}\exp(f_{q,p}(z))\pi_{0}(z)dz\approx\sum_{i=1}^{N_{\text{quad}}}w_{i},G(z_{i}^{\text{node}}),$ (6)
where $G(z)=\exp(f_{q,p}(z))\pi_{0}(z)$, and $\{z_{i}^{\text{node}},w_{i}\}_{i=1}^{N_{\text{quad}}}$ are the quadrature nodes and weights for the interval $\left[z_{\text{min}},z_{\text{max}}\right]$. The availability of $Z_{q,p}$ enables ITS, and the smoothness of $G$, together with its bounded support, makes this approximation accurate.
With this valid form for the normalized CDF, inverse transform sampling (ITS) can be used to draw samples from $p_{q,p}$ as measurable transformations of uniform noise. This preserves the deterministic structure of KART while enabling flexible, learnable latent distributions, well-founded so long as the KAN functions admit exponential integrability.
@@
As a result, the model retains a coherent probabilistic interpretation of KART that is consistent with the energy-based framework proposed by pang2020learninglatentspaceenergybased, which is well suited to maximum likelihood training. However, this approach is significantly limited, as it ignores inter-dimensional dependencies in the latent space and results in an axis-aligned joint distribution. We address this inflexibility in Sec. 2.3.2 by relaxing our adherence to KART.
2.1.1 Multivariate prior notation
From this point forward, we formulate our derivations over the multivariate probability space, $\bigl(\mathcal{Z},\mathcal{B}(\mathcal{Z}),\mathbb{P}_{\bm{\theta}}\bigr),$ where $\mathbb{P}_{\bm{\theta}}$ denotes a product measure parameterized by abstract model parameters, $\theta$, and the latent space is defined as a Cartesian product:
$\mathbb{P}_{\bm{\theta}}\;:=\;\bigotimes_{q=1}^{2n_{z}+1}\;\bigotimes_{p=1}^{n_{z}}\pi_{q,p},\quad\quad\quad\mathcal{Z}\;=\;\prod_{q=1}^{2n_{z}+1}\prod_{p=1}^{n_{z}}\mathcal{Z}_{j,k}\;\subset\;\mathbb{R}^{(2n_{z}+1)\times n_{z}},$ (7)
Each $\pi_{q,p}$ is a univariate probability measure on $\mathcal{Z}_{q,p}$ defined via exponential tilting, as introduced in Eq. 4. Under $\mathbb{P}_{\bm{\theta}}$, the latent variables $\{z_{q,p}\}_{q,p}$ are mutually independent, and the full latent variable
$\bm{z}\;=\;\bigl(z_{q,p}\bigr)_{q=1,\ldots,2n_{z}+1;\;p=1,\ldots,n_{z}}$
is a matrix-valued random variable whose entries correspond to independent univariate features.
2.2 Kolmogorov-Arnold Network generator
@@
Having sampled the intermediary latent variable with ITS such that $z_{q,p}=F^{-1}_{\pi_{q,p}}(u_{p})$, the generator can be realized flexibly as $\mathcal{G}:\mathcal{Z}\to\mathcal{X}$:
@@
$\bm{\tilde{x}}^{(s)}$ $=\mathcal{G}\left(\bm{z}^{(s)}\right)+\bm{\epsilon}^{(s)},\quad\text{where}\quad z_{q,p}^{(s)}\sim p_{q,p}(z),\quad\bm{\epsilon}^{(s)}\sim\mathcal{N}(\bm{0},\sigma_{\epsilon}^{2}\bm{I}),$ (8)
@@
where $\mathbf{\tilde{x}}^{(s)}=\{\tilde{x}_{o}^{(s)}\}_{o=1}^{n_{x}}$ represents a generated data sample corresponding to a single latent sample indexed by $s$, and $n_{x}$ is used to specify the dimension of $\bm{\tilde{x}}^{(s)}$. The additive noise is a Gaussian sample with hyperparameter $\sigma_{\epsilon}$. However, the architecture may be made strictly adherent to KART using a Kolmogorov-Arnold Network:
@@
$\tilde{x}_{o}^{(s)}$ $=\sum_{q=1}^{2n_{z}+1}\Phi_{o,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)+\epsilon_{o}^{(s)},\quad\mathbf{\tilde{x}}^{(s)}=\begin{pmatrix}\sum_{q=1}^{2n_{z}+1}\Phi_{1,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)\\[5.0pt] \sum_{q=1}^{2n_{z}+1}\Phi_{2,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)\\[5.0pt] \vdots\\[5.0pt] \sum_{q=1}^{2n_{z}+1}\Phi_{n_{x},q}\left(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\right)\end{pmatrix}+\bm{\epsilon}^{(s)}.$ (9)
The functions, $\Phi_{o,q}:\mathbb{R}\to\mathbb{R}$, are realizable as basis or KAN functions similar to $f_{q,p}:\mathcal{Z}_{q,p}\to\mathbb{R}$.
2.3 Scaling KAEM
The trainable components of the prior and inner functions are shared across the features of each input $\bm{x}^{(b)}$, while the outer generator functions remain separate for each output dimension. Specifically, the learnable functions are defined as:
$\bm{f}$ $=\begin{pmatrix}\left(f_{1,1},\dots,f_{1,n_{z}}\right)\\ \left(f_{2,1},\dots,f_{2,n_{z}}\right)\\ \vdots\\ \left(f_{2n_{z}+1,1},\dots,f_{2n_{z}+1,n_{z}}\right)\end{pmatrix},\quad\bm{\Phi}=\begin{pmatrix}\left(\Phi_{1,1},\dots,\Phi_{1,2n_{z}+1}\right)\\ \left(\Phi_{2,1},\dots,\Phi_{2,2n_{z}+1}\right)\\ \vdots\\ \left(\Phi_{n_{x},1},\dots,\Phi_{n_{x},2n_{z}+1}\right)\end{pmatrix}$, (10)
where $n_{x}$ denotes the dimensionality of the observed data.
2.3.1 Generator
The generator, $\mathcal{G}$, can be considered a separate generator network allowing for more flexible realizations, such as convolutional layers when adherence to Eq. 1 is not required. Deepening this generator, as guided by liu2024kankolmogorovarnoldnetworks for the Kolmogorov-Arnold Network (KAN), does not replace KART as the innermost foundation of KAEM:
$x_{o}^{(s)}=$ $\sum_{l_{n}=1}^{L_{n}}h_{o,l_{n}}\Bigg(\dots h_{l_{3},l_{2}}\Bigg(\sum_{l_{1}=1}^{L_{1}}h_{l_{2},l_{1}}\Bigg(\sum_{q=1}^{2n_{z}+1}\Phi_{l_{1},q}\Bigg(\sum_{p=1}^{n_{z}}z_{q,p}^{(s)}\Bigg)\Bigg)\Bigg)\Bigg)+\epsilon_{o}^{(s)}.$ (11)

### v9

@@ B.1 Inner functions as Markov kernels
In Eq. 13, $F^{-1}_{\pi_{q,p}}:[0,1]\to\mathcal{Z}$ is the generalized inverse CDF associated with a parameterized latent measure, $\pi_{q,p}$ in the latent space, $\mathcal{Z}\subset\mathbb{R}$. We have therefore interpreted $u_{p}\sim\text{Unif}(0,1)$ in Eq. 1 as a uniform random variable in the probability space, $\left([0,1],\mathcal{B}([0,1]),\mu\right)$, where $\mathcal{B}\left([0,1]\right)$ is the Borel $\sigma$-algebra on $[0,1]$ and $\mu$ is the uniform probability measure on $[0,1]$. This is valid under standard regularity assumptions on $F$, since it is a measurable, monotone, increasing function, and thus $F_{\pi_{q,p}}^{-1}(u)=\inf\left\{z\in\mathcal{Z}:F_{\pi_{q,p}}(z)\geq u\right\}$ is a measurable pushforward (Markov kernel) between the probability spaces:
$\left([0,1],\mathcal{B}\left([0,1]\right),\mu\right)\;\underset{F_{\pi_{q,p}}^{-1}}{\longrightarrow}\;\left(\mathcal{Z},\mathcal{B}\left(\mathcal{Z}\right),\pi_{q,p}\right),$ (26)
Under exponential tilting, the latent measure $\pi_{q,p}$ against a base measure $\pi_{0}$ can therefore be viewed from the lens of the Radon–Nikodym derivative:
@@
where $f_{q,p}(z)$ is our learned energy function and $Z_{q,p}$ is the finite normalizing constant (partition function) that ensures $\pi_{q,p}$ induces a valid probability density, $p_{q,p}$, for each latent feature. The inspiration for this exponential tilting form can be traced to the energy-based prior introduced by Pang et al. (2020), adapted to the univariate case. The base prior $\pi_{0}$ serves as a reference that guides the form of the learned $\pi_{q,p}$, allowing for flexible reweighting via the energy function $f_{q,p}$. For $\pi_{q,p}$ to define a valid density, which we denote $p_{q,p}$, we require exponential Lebesgue-integrability:
@@
$\text{such that}\quad Z_{q,p}$ $\;=\;\int_{\mathcal{Z}}\exp(f_{q,p}(z))\pi_{0}(dz)<\infty.$ (28)
@@
As a result, the model retains a coherent probabilistic interpretation of Eq. 1 that is consistent with the energy-based framework proposed by Pang et al. (2020), which is well suited to maximum likelihood training.
B.2 Kolmogorov-Arnold Network generator
@@
Having sampled the intermediary latent variable with ITS such that $z_{q,p}=F^{-1}_{\pi_{q,p}}(u_{p})$, the generator can be realized as:
@@
$\bm{\tilde{x}}$ $\;=\;G_{\Phi}(\bm{z})+\bm{\varepsilon},\quad\text{where}\quad\bm{z}\sim p_{f}(\bm{z}),\quad\bm{\varepsilon}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{noise}}\mathbf{I}),$ (29)
@@
where $\bm{\tilde{x}}=\{\tilde{x}_{o}\}_{o=1}^{n_{x}}$ represents a generated data sample and $n_{x}$ is used to specify the dimension of $\bm{\tilde{x}}$. The architecture may be made strictly adherent to KART using a Kolmogorov-Arnold Network (KAN):
@@
$\tilde{x}_{o}$ $=\sum_{q=1}^{2n_{z}+1}\Phi_{o,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}\right)+\varepsilon_{o},\quad\bm{\tilde{x}}=\begin{pmatrix}\sum_{q=1}^{2n_{z}+1}\Phi_{1,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}\right)\\[5.0pt] \sum_{q=1}^{2n_{z}+1}\Phi_{2,q}\left(\sum_{p=1}^{n_{z}}z_{q,p}\right)\\[5.0pt] \vdots\\[5.0pt] \sum_{q=1}^{2n_{z}+1}\Phi_{n_{x},q}\left(\sum_{p=1}^{n_{z}}z_{q,p}\right)\end{pmatrix}+\bm{\varepsilon}.$ (30)
The functions, $\Phi_{o,q}:\mathbb{R}\to\mathbb{R}$, are realizable as basis or KAN functions similar to $f_{q,p}:\mathcal{Z}\to\mathbb{R}$.
Deepening this generator, as guided by Liu et al. (2024) for KANs, does not replace KART as the inner foundation of KAEM:
$x_{o}=$ $\sum_{l_{n}=1}^{L_{n}}h_{o,l_{n}}\Bigg(\dots h_{l_{3},l_{2}}\Bigg(\sum_{l_{1}=1}^{L_{1}}h_{l_{2},l_{1}}\Bigg(\sum_{q=1}^{2n_{z}+1}\Phi_{l_{1},q}\Bigg(\sum_{p=1}^{n_{z}}z_{q,p}\Bigg)\Bigg)\Bigg)\Bigg)+\varepsilon_{o}.$ (31)